Jill has $3.50 in nickels and dimes. If she has 50 coins, how
many of each type of coin does she have?

Answer by checkley75(3666)

N+D=50 OR N=50-D
5N+10D=350
5(50-D)+10D=350
250-5D+10D=350
5D=350-250
5D=100
D=100/5
D=20 NUMBER OF DIMES.
50-20=30 NUMBER OF NICKLES.
PROOF
5*30+10*20=350
150+200=350
350=350